OpenStudy (anonymous):

How do you determine the domain and range of a function expressed with an equation? for example: y=2- sqrt X

Mehek (mehek14):

domain = any value you would put for x range = the output or the result of what you put for x

OpenStudy (anonymous):

how can i solve that equation?

OpenStudy (michele_laino):

I think that 1) the domain of a function is given by all values of the variable \(x\), such that the equation which expresses the function has a meaning whereas, 2) the range of a function, is given by the set of all values which gives the corresponding equation, when I replace \(x\) with any value of the range of such function

OpenStudy (michele_laino):

is your function like this: \(\Large y=2- \sqrt{x}\)

OpenStudy (anonymous):

yes :)

OpenStudy (michele_laino):

then we can say, that such equation has a meaning, if \(x\) is any non negative number, namely the domain of such function is the subsequent set: \[domain = \left\{ {x \in \mathbb{R},\quad {\text{such that }}x \geqslant 0} \right\}\] since there is not the square root of negative numbers

OpenStudy (anonymous):

thanks so much !

OpenStudy (michele_laino):

:)